Four added to the product of 3 and -11 is what number?
Please show me the steps to get the answer. Thanks

Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

What number results if 11 is subtracted from the product of -6 and 5?
-----
Ans: (-6)(5)-11 = -30-11 = -41
-------
Cheers,
Stan H.

Answer by Alan3354(69443)

Four added to the product of 3 and -11 is what number?
-------------
3*-11 + 4 = -29